OpenStudy (anonymous):

For hydrocyanic acid, HCN, Ka = 4.9 × 10-10. Calculate the pH of 0.20 M NaCN. What is the concentration of HCN in the solution? for the ph i got 10 is that right?

OpenStudy (aaronq):

What did you do?

OpenStudy (anonymous):

i did 4.9^-10 multiplied it times 0.20 and then i took the square root

OpenStudy (kash_thesmartguy):

pH = 8.63

OpenStudy (kash_thesmartguy):

not sure!

OpenStudy (aaronq):

You have to always look at what you're given. They gave the pKa for HCN but the solution used was NaCN, the conjugate base of HCN. The solution is acting as a base: \(\sf NaCN+H_2O\rightleftharpoons NaOH+HCN\) and you use the Kb, \(\sf K_B*K_A=K_W\) The expression is: \(\sf K_B=\dfrac{[NaOH][HCN]}{NaCN}\) By using I.C.E. table you arrive at: \(\sf [NaOH]=\sqrt{K_B*[NaCN]}\) Find pOH, then pH (note: pH+pOH=14) Also, \(\sf [HCN]=[NaOH]\)